OpenStudy (anonymous):

Factor the polynomial completely. (If the expression is not factorable using integers, enter PRIME.) 9x^2 − 4

OpenStudy (asnaseer):

can you spot that this is of the form:\[a^2-b^2\]

OpenStudy (anonymous):

its not prime for sure

OpenStudy (asnaseer):

do you know how to factorise expressions of this type?\[a^2-b^2\]

OpenStudy (anonymous):

its called the difference of squares isnt it?

OpenStudy (asnaseer):

yes - and it has a well known factorisation:\[a^2-b^2=(a+b)(a-b)\]in your problem: a=3x b=2

OpenStudy (anonymous):

(3x+2)(3x-2)

OpenStudy (asnaseer):

that is correct - well done!